# Probing LLMs for Multilingual Discourse Generalization Through a Unified Label Set

### v1

@@ 1 Introduction
We find that overall LLMs are able to generalize at the discourse level across languages and frameworks, and that multilingual training and larger model sizes both increase probe performance. Our layer-wise analyses show that language generalization at the discourse level is most salient in the intermediate layers, which are most predictive of multilingual discourse information. To our best knowledge, this is the first work to apply a unified label set to multilingual discourse relation classification at scale. Lastly, we discuss challenges and biases in LLM discourse representations, providing insights into the limitations and potential avenues for improving discourse modeling in multilingual and generalization settings.
@@ 2 Related Work
A similar effort to unify discourse relations has been made by the 2021 DISRPT Shared Task Zeldes et al. (2021) which introduced the first iteration of the discourse relation classification task in a unified format. It leveraged shared foundational assumptions across frameworks. However, no unified discourse relation labels were proposed, meaning that each dataset has its own label set, even for the ones that come from the same framework. To address this, we first conduct an extensive review of previous work on discourse relation mapping proposals and build upon existing work to present a unified label set, which we describe in Section 3.
@@
Previous work has examined the ability of PLMs/LLMs to understand discourse such as coreference resolution Gan et al. (2024); Saputa et al. (2024), but their investigations are either limited to framework-dependent representations, monolingual datasets, or focus on single domains. Specifically, Koto et al. (2021) examined a variety of PLMs through a set of framework-dependent probing tasks for discourse coherence by looking at the residual stream (i.e. token representations), while we approach discourse relation classification with a unified format and label set across various frameworks and $13$ languages (§3) using attentions (§4), offering opportunities for investigating discourse generalization across languages and frameworks. Kurfalı and Östling (2021) extended discourse probing to multilingual PLMs such as the multilingual BERT and XLM-RoBERTa (Devlin et al., 2019; Conneau et al., 2020) to examine how well they transfer discourse-level knowledge across languages, but their evaluation of discourse coherence was also framework-dependent and was only performed on two English datasets in the news domain. Lastly, Kim and Schuster (2023) studied discourse understanding in LMs by probing their ability to track discourse entities, but their investigation is also limited to English.
@@ 3 A Unified Label Set
Building up on previous work and effort on mapping discourse relations across corpora and frameworks Benamara and Taboada (2015); Bunt and Prasad (2016); Liu and Zeldes (2023), we present a unified set of $17$ discourse relation classes that are identified as core discourse relations by Bunt and Prasad (2016) and are attested in various domains and text types. This unified label set is motivated by both theoretical groundings and empirical studies, and takes annotation guidelines into considerations. Specifically, the proposed unified label set adapts the top-level classes from the mapping proposal described in Benamara and Taboada (2015) and extends it to phenomena frequent in dialogues such as acknowledgment, interruption, and correction Asher et al. (2016). We describe the top-level classes below and include definitions and examples for all $17$ relation labels in Appendix A.
temporal is mapped to framework-specific labels that establish a chronological sequence between events or states. Temporal relations indicate when one event occurs in relation to another such as before, after, or simultaneously. These relations help organize discourse by providing a timeline of events. RST’s sequence, PDTB’s Temporal.Asynchronous/Synchronous, and SDRT’s temploc and flashback (following Muller et al. (2012)) all fall under this class.
@@
causal-argumentative contains subclasses that indicate a causal relation or involve rhetorical reasoning that shapes the coherence and persuasiveness of an argument: causal, adversative, explanation, evaluation, contingency, and enablement. adversative is defined as connecting discourse units for which some incompatibility is being highlighted and covers commonly used discourse relations in virtually all frameworks such as concession and contrast.
@@ 4 Probing LLMs for Discourse Relations
In the computational graph of the Transformer, attention layers provide the only connections between the next token prediction and previous token positions of the input sequence. We therefore hypothesize that generating predictions relying on discourse-level information necessarily has to involve the connections provided by the self-attention layers. To uncover the extent to which discourse information is represented, we thus propose to probe the attention scores between the tokens contained in the two input sequences for discourse relation classification.
Probing requires a fixed length representation to enable training a classifier. Inspired by Alain and Bengio (2018) and Koto et al. (2021), we use maximum pooling to turn the attention scores of variable token-length sentences into a fixed-length relational representation. To be more precise, for a document $d$ of length $N$ tokens, we propose to compute the full attention matrix $\mathbf{X}\in\mathbb{R}^{L\times H\times N\times N}$ where $L$ is the number of layers and $H$ is the number of attention heads in each layer. Let $I_{1}=(i_{1},i_{2},...,i_{N_{1}})$, $I_{2}=(j_{1},j_{2},...,j_{N_{2}})$ be the token indices of two sequences $s_{1},s_{2}\subset d$ where $I_{1}<I_{2}$. The pooling step is carried out for each attention head in each layer. We thus have:
@@
In other words, $\mathbf{C},\mathbf{D}_{1},\mathbf{D}_{2}$ are the matrices of pooled attention scores between and within the two spans respectively (we only consider the lower half of the attention matrix, as for decoder models the upper half is masked). We probe on the flattened concatenation $\mathbf{a}\in\mathbb{R}^{3LH}$ of these representations. Here, we follow Tenney et al. (2019) and use a two-layer MLP with $\mathrm{tanh}$ and Sigmoid activations $\sigma$. Our probe is thus a classifier of the form
@@
We probe both the combined attention scores over all layers as well as the layer-wise representations. [^1] https://github.com/mainlp/discourse_probes We hypothesize that the former will lead to better overall probe accuracy as the understanding of more complex discourse relations typically requires hierarchical processing steps. We include the latter layer-wise analysis to study the structural dynamics of discourse processing in the model.
@@ 6.1 Overall Performance and Comparison
Figure 2 shows the performance of our probes on DISRPT using the unified label set for the $23$ LLMs along with their sizes. [^3] See Table 2 in Appendix C for a detailed overview. Overall, a clear trend emerges wherein larger models generally achieve higher accuracy, but with notable deviations based on their language coverage. Interestingly, even the Llama3-3B probe is able to outperform the fine-tuned DisCoDisCo trained on all languages ($49.1$% vs $47.9$%), while the larger model probes beat it by a margin of up to $10.3$%. Despite their multilingual training, the BLOOM model probes underperform compared to other LLMs and DisCoDisCo. Several factors may be attributed to this including smaller training data, tokenizer effect Toraman et al. (2023), and limited coverage of languages present in DISRPT ($54$%: French, Spanish, Portuguese, English, and Chinese). This aligns with and supplements Dakle et al. (2023)’s findings of the evaluation of BLOOM models on a variety of syntactic and semantic tasks as well as their unsatisfying performance in multilingual settings. In addition, Llama3 exhibits the lowest proportion of DISRPT languages covered, which does not seem to lead to a distinct disadvantage as its probes outperform the similarly sized smaller versions of Qwen2.5 and match its larger counterparts, reporting a higher coverage of ‘supported’ languages.
@@ 6.2 Language Generalization
There is one exception to this generalization trend: for the Thai dataset, the mono-probe outperforms multi-all-probe with a margin of $3.2\%$. This is likely due to the fact that the Thai dataset is relatively large and contains only news data and was only annotated with explicit discourse relations Prasertsom et al. (2024), which has been reported to be considerably easier than implicit discourse relations Knaebel (2021); Braud et al. (2023) due to the presence of connectives, which, while not unambiguous, help narrow down the likely senses of relations Webber et al. (2019). The probes for Basque consistently underperform, with the multi-all-probe achieving the same performance as the Basque-only probe. This is likely due to the fact that Basque is considered a language isolate, a language that has no demonstrable genetic relationship with any other language Campbell (2010). Besides, most models do not include it in their list of supported languages (see Table 2 in Appendix). The reference system DisCoDisCo also does not generalize well for Basque: its accuracy drops by $26.4$% compared to the dataset-specific model. Moreover, the mono-probes for Basque and Turkish exhibit lower accuracy, which might mean that models trained on discourse tasks need more language-specific adaptations, particularly for morphologically rich and typologically different languages.
@@
Training on discourse relation instances from related languages often leads to better generalization than training on a single language. For instance, the multi-lang-probe for the Indo-European languages achieves reasonable performance across the Romance languages in DISRPT: there are improvements over the mono-probe for Portuguese ($+2.5$%), Spanish ($+6.5$%), and Italian ($+3.1$%). For the Germanic and Romance language groups, we observe a similar generalization effect, though to a lesser extent. Notably, the Germanic multi-lang-probe leads to the highest accuracy on the German test set ($+0.8$% over the mono-probe). These are encouraging results as some of these languages do not have a large amount of instances in DISRPT: from the total number of data, Spanish covers $1.7$%, and Italian only $0.7$% (see Table 1 in Appendix B). This suggests that leveraging discourse relation annotations from related language as well as the same underlying framework help with generalization. Adding to that, training on English only, which has the largest number of samples covering more than half of DISRPT, shows surprising generalization effects to other languages, with Portuguese and Spanish achieving an accuracy of $48.2\%$ and $49\%$ respectively, and French reaching $45.3\%$.
@@ 6.3 Layer-wise Analysis
The layer-wise probes reveal additional insights into the structural processing of discourse information within the model. Firstly, Figure 4a shows the test set performance by language families. Here, we observe a trend by which the probes in lower to middle layers are the most performant as accuracy improves rapidly in the early layers and stabilizes around layer 1010101015151515. In particular, for Turkish and Thai, the probe accuracy drops by almost $50$% in higher layers. This effect is also observed for the Chinese probe but to a smaller extent. This shows that the discourse representations are best aligned across languages in the middle layers, which could be explained by the findings of Wendler et al. (2024) according to which multilingual computations are carried out in an English-aligned “concept space” in the intermediate layers. Surprisingly, the final layer probes show another increase for Chinese, Turkish, and Thai, which is also reflected in the overall test set accuracy.
@@ 6.4 Error Analysis
To better understand the probe performance, we plot a confusion matrix in Figure 5 to study systematic errors. Overall, elaboration is the most frequent label. Here, our probe has a bias of confusing it with labels such as framing and explanation. This suggests that the model struggles to differentiate between content expansion (elaboration) which sometimes overlaps semantically with explanation, and relations that involve additional information but with a primary focus on contextualization (framing). In other words, elaboration is mainly used to give additional information about an entity or a proposition, while framing provides information with a goal to increase a reader’s understanding of an entity or a proposition Mann and Thompson (1988); Carlson and Marcu (2001). This corresponds to the distinction made by Hovy and Maier (1997): elaboration is considered an ideational class that is used to express states of affairs in the world, not including the interlocutors; on the other hand, framing and explanation are considered interpersonal and are expressed by various perlocutionary acts to affect readers’ attitude and beliefs. elaboration is also a highly prevalent relation in general, making it a probable default prediction in ambiguous cases.
@@ 7 Conclusion
Our probes exhibit generalization, whereby training across languages generally outperforms language–specific probes, which is not the case for the reference system DisCoDisCo. Through multi-faceted analyses, we find that model size alone does not lead to discourse probing success; instead, multilingual training, dataset composition, and language-specific factors play significant roles. While larger models generally perform better, discrepancies such as the under-performing BLOOM and the best-performing Aya-23-35B-probe emphasize the importance of training data quality and architectural optimizations. Surprisingly, our English mono-probe generalizes to Romance languages. Our layer-wise analysis suggests that discourse representations are best aligned across languages in the intermediate layers, with later layers refining these representations for specific relation types. In addition, models struggle with relations requiring implicit reasoning such as explanation.
Overall, our findings highlight the interplay between multilinguality, scaling, and internal representations of LLMs for multilingual discourse processing and provide insights into cross-lingual alignment of discourse relations. Future work can improve cross-lingual alignment by refining representations of challenging relation classes. By advancing our understanding of discourse generalization in LLMs, we contribute towards more interpretable and robust NLP systems capable of nuanced language comprehension.
@@
Secondly, our approach assumes a unified label set for discourse relations across languages and frameworks, which we present in Section 3. While this enables cross-linguistic and cross-framework discourse analysis, compromises were necessary, which to some extent simplify the complexity and granularity of discourse relations assumed by frameworks such as PDTB and RST. Fully standardizing and harmonizing discourse relations is inherently challenging, and finding a good trade-off between maintaining theoretical assumptions and ensuring practical applicability is crucial yet complex, as evidenced in previous mapping proposals and efforts Benamara and Taboada (2015); Demberg et al. (2019). In addition, segmentation differences exist across frameworks, which can have an impact on the performance of our probes for certain relations such as attribution. This work should thus be interpreted with an awareness of the theoretical and practical difficulties in creating an informed and unified taxonomy suitable for both theoretical studies and computational research. This work should also be viewed as facilitating the development of a unified discourse relation representation for computational discourse modeling such as the effort and initiatives that have been put forward for dependency parsing (Universal Dependencies, de Marneffe et al. 2021), empirical study of anaphora (Universal Anaphora, Poesio et al. 2024), and semantic parsing (Uniform Meaning Representation, Bonn et al. 2024).

### v2

@@ 1 Introduction
We find that overall LLMs are able to generalize at the discourse level across languages and frameworks, and that multilingual training and larger model sizes both increase probe performance. Our layer-wise analyses show that language generalization at the discourse level is most salient in the intermediate layers, which are most predictive of multilingual discourse information. To our best knowledge, this is the first work to apply a unified label set to multilingual discourse relation classification at scale. Lastly, we discuss challenges and biases in LLM discourse representations, providing insights into the limitations and potential avenues for improving discourse modeling in multilingual and generalization settings. The implementations of our experiments are available on GitHub. [^1] https://github.com/mainlp/discourse_probes
@@ 2 Related Work
In particular, while Bunt and Prasad (2016) identified a set of core discourse relations, it did not cover DEP and was limited to English and French only. The examined corpora in their work also did not cover discourse phenomena concerning pragmatics or textual organization, both of which are indispensable aspects in discourse analysis. For instance, background and motivation are two RST-style relations that are not present in the examined RST Discourse Treebank (RST-DT, Carlson et al. 2003). Both relations are expressed by various perlocutionary acts to affect readers’ or speakers’ attitude and beliefs. To address these limitations, we first conduct an extensive review of previous work on discourse relation mapping proposals and present a unified label set (§3) to enable empirical studies in a multilingual setting, which has not been systematically explored before.
The 2021 DISRPT Shared Task Zeldes et al. (2021) introduced the first iteration of the discourse relation classification task in a unified format. It leveraged shared foundational assumptions across frameworks. However, no unified discourse relation labels were proposed, meaning that each dataset has its own label set, even for the ones that come from the same framework. We thus leverage this resource and propose a unified label set that is the first to be empirically tested in discourse relation classification across $13$ languages, four frameworks, and $26$ datasets, which cover various, modern genres, domains, and modalities.
@@
Previous work has examined the ability of PLMs/LLMs to understand discourse Gan et al. (2024); Saputa et al. (2024); Miao et al. (2024), but their investigations are either limited to framework-dependent representations, monolingual datasets, or focus on single domains. Specifically, Koto et al. (2021) examined a variety of PLMs through a set of framework-dependent probing tasks for discourse coherence by looking at the residual stream (i.e. token representations), while we approach discourse relation classification with a unified format and label set across various frameworks and $13$ languages (§3) using attentions (§4), offering opportunities for investigating discourse generalization across languages and frameworks. Kurfalı and Östling (2021) extended discourse probing to multilingual PLMs such as the multilingual BERT and XLM-RoBERTa (Devlin et al., 2019; Conneau et al., 2020) to examine how well they transfer discourse-level knowledge across languages, but their evaluation of discourse coherence was also framework-dependent and was only performed on two English datasets in the news domain. Lastly, Kim and Schuster (2023) studied discourse understanding in LMs by probing their ability to track discourse entities, but their investigation is also limited to English.
@@ 3 A Unified Label Set
Building up on previous effort on mapping discourse relations across corpora and frameworks Benamara and Taboada (2015); Bunt and Prasad (2016); Liu and Zeldes (2023), we present a unified set of $17$ discourse relation classes to facilitate empirical investigation that is not constrained by framework-dependent discourse representations. This unified label set is motivated by both theoretical groundings and empirical studies, and takes annotation guidelines into considerations. Specifically, the proposed unified label set adapts the top-level classes from the mapping proposal described in Benamara and Taboada (2015) and extends it to phenomena frequent in dialogues such as acknowledgment, interruption, and correction Asher et al. (2016). Through a series of experiments, we demonstrate how our unified label set generalizes across languages and frameworks, providing a foundation for future empirical studies. Below we describe the top-level classes and include definitions and examples for all $17$ relation labels in Appendix A.
temporal is mapped to framework-specific labels that establish a chronological sequence between events or states. Temporal relations indicate when one event occurs in relation to another such as before, after, or simultaneously. These relations help organize discourse by providing a timeline of events. RST’s sequence, PDTB’s Temporal.Asynchronous/Synchronous, and SDRT’s temploc and flashback (following Muller et al. 2012) all fall under this class.
@@
Figure 2: Schematic visualization of how the attention representations are obtained. We probe the concatenation of pooled representations $(\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C})$. Scores mapping from or to context-tokens are not considered. Note that we study decoder-only models where attention maps only to previous tokens.
causal-argumentative contains subclasses that indicate a causal relation or involve rhetorical reasoning that shapes the coherence and persuasiveness of an argument: causal, adversative, explanation, evaluation, contingency, and enablement. adversative is defined as connecting discourse units for which some incompatibility is being highlighted and covers commonly used discourse relations such as concession and contrast.
@@ 4 Probing LLMs for Discourse Relations
In the computational graph of the Transformer, attention layers provide the only connections between the next token prediction and previous token positions of the input sequence. We therefore argue that generating predictions relying on discourse-level information necessarily has to involve the connections provided by the self-attention layers. To uncover the extent to which discourse information is represented, we thus propose to probe the attention scores between the tokens contained in the two input sequences for discourse relation classification. Note that we do not assume that the processing of discourse-level phenomena is exclusively located in the attention layers. Rather, observing attention scores provides a lightweight and scalable approach to our research question.
Probing requires a fixed length representation to enable training a classifier. Inspired by Alain and Bengio (2018) and Koto et al. (2021), we use maximum pooling to turn the attention scores of variable token-length sentences into a fixed-length relational representation. Our approach is illustrated in Figure 2. To be more precise, for a document $d$ of length $N$ tokens, we propose to compute the full attention matrix $\mathbf{X}\in\mathbb{R}^{L\times H\times N\times N}$ of attention scores where $L$ is the number of layers and $H$ is the number of attention heads in each layer. We do this by inputting the whole document $d$ into the model at once computing a single forward pass. Since we are only interested in the model internals, we ignore the outputs and thus do not impose any specific decoding strategy. Let $I_{1}=(i_{1},i_{2},...,i_{N_{1}})$, $I_{2}=(j_{1},j_{2},...,j_{N_{2}})$ be the token indices of two sequences $s_{1},s_{2}\subset d$ where $I_{1}<I_{2}$. The pooling step is carried out for each attention head in each layer. We thus have:
@@
In other words, $\mathbf{C},\mathbf{D}_{1},\mathbf{D}_{2}$ are the matrices of maximum-pooled attention scores between and within the two spans respectively (we only consider the lower half of the attention matrix, as for decoder models the upper half is masked). We also ablate other strategies such as mean pooling and using subsets of the attention scores described (see Table 4 and Table 5 in Appendix D), and find that the proposed setup strikes the best balance between performance and size of the representations. We probe the flattened concatenation
$\mathbf{a}=flatten((\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}))\in\mathbb{R}^{3LH}$ (2)
Here, we follow Tenney et al. (2019) and use a two-layer MLP with $\mathrm{tanh}$ and Sigmoid activations $\sigma$. Our probe is thus a classifier of the form
@@
We probe both the combined attention scores over all layers as well as the layer-wise representations. We hypothesize that the former will lead to better overall probe accuracy as the understanding of more complex discourse relations typically requires hierarchical processing steps. We include the latter layer-wise analysis to study the structural dynamics of discourse processing in the model.
@@ 6.1 Overall Performance and Comparison
Figure 3 shows the performance of our probes on DISRPT using the unified label set for the $23$ LLMs along with their sizes (see Table 2 in Appendix C for a detailed overview). Overall, a clear trend emerges wherein larger models generally achieve higher accuracy, but with notable deviations based on their language coverage. Interestingly, even the Llama3-3B probe is able to outperform the fine-tuned DisCoDisCo trained on all languages ($49.1$% vs $47.9$%), while the larger model probes beat it by a margin of up to $10.3$%. Despite their multilingual training, the BLOOM model probes underperform compared to other LLMs and DisCoDisCo. Several factors may be attributed to this including smaller training data, tokenizer effect Toraman et al. (2023), and limited coverage of languages present in DISRPT ($54$%: French, Spanish, Portuguese, English, and Chinese). This is consistent and further supplements Dakle et al. (2023)’s findings on the evaluation of BLOOM models across a variety of syntactic and semantic tasks as well as their unsatisfying performance in multilingual settings. In addition, Llama3 exhibits the lowest proportion of DISRPT languages covered, which does not seem to lead to a disadvantage as its probes outperform the similarly sized smaller versions of Qwen2.5 and match its larger counterparts, reporting a higher coverage of ‘supported’ languages.
@@ 6.2 Language Generalization
There is one exception to this generalization trend: for the Thai dataset, the mono-probe outperforms multi-all-probe with a margin of $3.2\%$. This is likely due to the fact that the Thai dataset is relatively large and contains only news data and was only annotated with explicit discourse relations Prasertsom et al. (2024), which has been reported to be considerably easier than implicit discourse relations Knaebel (2021); Braud et al. (2023) due to the presence of connectives, which, while not unambiguous, help narrow down the likely senses of relations Webber et al. (2019). The probes for Basque consistently underperform, with the multi-all-probe achieving the same performance as the Basque-only probe. This is likely due to the fact that Basque is considered a language isolate, a language that has no demonstrable genetic relationship with any other language Campbell (2010). Besides, most models do not include it in their list of supported languages (see Table 2 in Appendix C). The reference system DisCoDisCo also does not generalize well for Basque: its accuracy drops by $26.4$% compared to the dataset-specific model. Moreover, the mono-probes for Basque and Turkish exhibit lower accuracy, which might mean that models trained on discourse tasks need more language-specific adaptations, particularly for morphologically rich and typologically different languages.
@@
Training on discourse relation instances from related languages often leads to better generalization than training on a single language. For instance, the multi-lang-probe for the Indo-European languages achieves reasonable performance across the Romance languages in DISRPT: there are improvements over the mono-probe for Portuguese ($+2.5$%), Spanish ($+6.5$%), and Italian ($+3.1$%). For the Germanic and Romance language groups, we observe a similar generalization effect, though to a lesser extent. Notably, the Germanic multi-lang-probe leads to the highest accuracy on the German test set ($+0.8$% over the mono-probe). These are encouraging results as some of these languages do not have a large amount of instances in DISRPT: from the total number of data, Spanish covers $1.7$%, German $1.19$%, and Italian only $0.7$% (see Table 1 in Appendix B). This suggests that leveraging discourse relation annotations from related language as well as the same underlying framework help with generalization. Adding to that, training on English only, which has the largest number of samples covering more than half of DISRPT, shows surprising generalization to other languages, with Portuguese and Spanish achieving an accuracy of $48.2\%$ and $49\%$ respectively, and French reaching $45.3\%$.
@@ 6.3 Layer-wise Analysis
The layer-wise probes reveal additional insights into the structural processing of discourse information within the model. Firstly, Figure 5a shows the test set performance by language families. Here, we observe a trend by which the probes in lower to middle layers are the most performant as accuracy improves rapidly in the early layers and stabilizes around layer 1010101015151515. In particular, for Turkish and Thai, the probe accuracy drops by almost $50$% in higher layers. This effect is also observed for the Chinese probe but to a smaller extent. This shows that the discourse representations are best aligned across languages in the middle layers, which could be explained by the findings of Wendler et al. (2024) according to which multilingual computations are carried out in an English-aligned “concept space” in the intermediate layers. This result is also aligned with concurrent work by Skean et al. (2025) which finds that final-layer representations are consistently outperformed by intermediate representations across a range of tasks and architectures. Our work supplements this by extending their finding to linguistic tasks such as discourse relation classification.
Surprisingly, the final layer probes show another increase for Chinese, Turkish, and Thai, which is also reflected in the overall test set accuracy. This could indicate that their probes benefit from language-specific features, which might be more prevalent in layers close to the outputs if we assume a shared, compressed representation space in intermediate layers as proposed by Wendler et al. (2024) and Deletang et al. (2024).
@@ 6.4 Error Analysis
To better understand the probe performance, we plot a confusion matrix in Figure 6 to study systematic errors. Overall, elaboration is the most frequent label. Here, our probe has a bias of confusing it with labels such as framing and explanation. This suggests that the model struggles to differentiate between content expansion (elaboration) which sometimes overlaps semantically with explanation, and relations that involve additional information but with a primary focus on contextualization (framing).
@@
In other words, elaboration is mainly used to give additional information about an entity or a proposition, while framing provides information with a goal to increase a reader’s understanding of an entity or a proposition Mann and Thompson (1988); Carlson and Marcu (2001). This corresponds to the distinction made by Hovy and Maier (1997): elaboration is considered an ideational class that is used to express states of affairs in the world, not including the interlocutors; on the other hand, framing and explanation are considered interpersonal and are expressed by various perlocutionary acts to affect readers’ attitude and beliefs. elaboration is also a highly prevalent relation in general, making it a probable default prediction in ambiguous cases.
@@ 7 Conclusion
Our probes exhibit generalization, whereby training across languages generally outperforms language–specific probes, which is not the case for the reference system DisCoDisCo. Through multi-faceted analyses, we find that model size alone does not lead to discourse probing success; instead, multilingual training, dataset composition, and language-specific factors play significant roles. While larger models generally perform better, discrepancies such as the under-performing BLOOM and the best-performing Aya-23-35B-probe emphasize the importance of training data quality and architectural optimizations. Surprisingly, we find that some of our probes generalize to languages unseen during probe training. For example, the English mono-probe generalizes to Romance languages. Furthermore, our layer-wise analysis suggests that discourse representations are best aligned across languages in the intermediate layers, with later layers refining these representations for specific relation types. In addition, models struggle with relations requiring implicit reasoning such as explanation.
Overall, our findings highlight the interplay between multilinguality, scaling, and internal representations of LLMs for multilingual discourse processing and provide insights into cross-lingual alignment of discourse relations. Understanding which discourse relations are well-captured by LLMs and which are not could help improve discourse parsing models by highlighting gaps in current representation. In addition, our findings may be useful for designing better discourse-aware pre-training or fine-tuning strategies. Future work can improve cross-lingual alignment by refining representations of challenging relation classes. More generally, our method provides a systematic way to investigate discourse representation encoded in LLMs, making it a useful tool for answering linguistic questions that can be formulated into a consecutive span representation. By advancing our understanding of discourse generalization in LLMs, we contribute towards more interpretable and robust NLP systems capable of nuanced language comprehension.
@@
Secondly, our approach assumes a unified label set for discourse relations across languages and frameworks, which we present in §3. While this enables cross-linguistic and cross-framework discourse analysis, compromises were necessary, which to some extent simplify the complexity and granularity of discourse relations assumed by frameworks such as PDTB and RST. Fully standardizing and harmonizing discourse relations is inherently challenging, and finding a good trade-off between maintaining theoretical assumptions and ensuring practical applicability is crucial yet complex, as evidenced in previous mapping proposals and efforts Benamara and Taboada (2015); Demberg et al. (2019). In addition, segmentation differences exist across frameworks, which can have an impact on the performance of our probes for certain relations such as attribution. This work should thus be interpreted with an awareness of the theoretical and practical difficulties in creating an informed and unified taxonomy suitable for both theoretical studies and computational research. This work should also be viewed as facilitating the development of a unified discourse relation representation for computational discourse modeling such as the effort and initiatives that have been put forward for dependency parsing (Universal Dependencies, de Marneffe et al. 2021), empirical study of anaphora (Universal Anaphora, Poesio et al. 2024), and semantic parsing (Uniform Meaning Representation, Bonn et al. 2024).